Set up a system of equations and solve. Mike sold 12 sports cards. He sold baseball cards for 25 cents each and football cards for 5 cents each. Mike made a total of $2.00. How many of each type of card did Mike
sell?
My teacher just doesn't show us how to do word problems

Answer by Boreal(15235)

baseball cards=b
football cards=f
b+f=12, since he sold 12 cards
so let f=12-b
now there is only one variable to solve for
.25*b+.05(12-b)=2.00, the money equation
.25b+.60-0.05b=2.00
.20b=1.40
b=7baseball cards of 0.25 or $1.75
12-b=5 football cards of 0.05 or $0.25
They add to $2.00